# Class Agnostic Instance-level Descriptor for Visual Instance Search

### v1

@@ III-B Decomposition Seed Selection
$A_{ij}=\begin{cases}1,&x_{i}^{\prime}{\cdot}x_{j}>\alpha,\\ 0,&otherwise,\\ \end{cases}$ (5)
@@
[…]
So $x_{b}$ and $x_{w}$ are selected as two seeds for the bisecting. The rest of features in $S$ are assigned to cluster of $x_{b}$ or $x_{w}$ by checking whether they are closer to $x_{b}$ or $x_{w}$. The initialization scheme is adopted for the bisecting of all the nodes on the hierarchy. The difference between the clustering with seed selection and clustering with random initialization is illustrated in Figure 2.
@@ III-C Terminate Condition
Given a feature set (or a node) $S$ on the hierarchy, the bisecting over it should not be carried out when the connectivity within the feature set is already sufficiently tight. The integral of these features inside usually represents a compact semantic region in the image. Given the total number of edges within the feature set $S$ is
$c(S)=\sum_{i,j\in S\&i<j}A_{ij},$ (8)
@@
where $n=|S|$. Apparently, $\bar{c}(S)$ is in the range $[0,1]$. During the top-down decomposition, as the bisecting is carried out on a deeper node, the feature $S$ becomes more compact. $\bar{c}(S)$, therefore, gets higher. We can control decomposition granularity by tuning the threshold $\tau_{1}$ on $\bar{c}(S)$. A larger threshold leads to finer decomposition granularity. In our implementation, when $\bar{c}(S)\geq{\tau_{1}}$, the bisecting on $S$ is not undertaken further across all the experiments.
@@ III-D Condition on Dummy Nodes
Given a feature set (or a node) $S$ on the hierarchy, it does not necessarily correspond to a semantically compact region. It could be a mixture of multiple instances of different categories or the mixture of objects and the background, or simply empty background. Such feature set is called “dummy node”. It is not meaningful to build instance-level feature from it. Inspired by [32], “dummy node” is judged by the strength of feature energy. In our method, we define $H=\arg\max\nolimits_{\forall x_{i}\in X}^{\theta}\left|x_{i}\right|^{1}$ as the set of features that exhibit high energy within the dataset $X$. Here, we set $\theta$ to the top-$30\%$ based on the relevant experimental results presented in [32]. We measure the overlapping ratio between $S$ and $H$
@@
When $\xi(S)$ is low, it basically indicates the region is over-dominated by semantically insignificant features. A threshold $\tau_{2}$ is set to ignore these dummy nodes on the hierarchy. No feature will be built from it when a detected region is viewed as “dummy node”. However, this condition does not prevent the feature set $S$ from further bisecting. It is possible that the direct/indirect subset of $S$ becomes a salient region when the bisecting reaches a more fine-grained level.
Figure 3: The framework of building instance-level feature for instance search. There are two major components, instance-level region detector and feature descriptor. Given an image, a set of patch-level features are produced by the self-supervised backbone. The hierarchical decomposition is applied on the feature set. Each node on the hierarchy corresponds to a potential instance region or sub-region. The patch-level masks can be produced for the valid nodes on the hierarchy. The feature for each region is pooled from another network. Given an incoming query, feature is extracted with the same backbone.
Input: Input image $I$
@@ 2 Initialize $\mathcal{L}\leftarrow X$ and $\mathcal{R}\leftarrow\emptyset$;
4      Pop $node$ from $\mathcal{L}$;
@@
6            $\mathcal{R}$.append($node$);
@@ 10             $\mathcal{L}\leftarrow\mathcal{L}\cup\text{getObject}(S_{b})\cup\text{getObject}(S_{w})$;
The entire process of instance-level detection is summarized in Algorithm 1. The hierarchical decomposition starts from feature set $X$. The bisecting is repeatedly applied on $X$ and its subsets. The resulting subsets $S_{b}$ and $S_{w}$ do not necessarily correspond to two instance regions. Several instances from the same object category are likely partitioned into one subset. As a result, one more step “$getObject(\cdot)$” (Algorithm 1 Line – 10) is undertaken to partition subsets $S_{b}$ and $S_{w}$ respectively into spatially connected smaller subsets. Two subsidiary conditions are involved in the algorithm. Specifically, if a node is dummy according to $\xi(node)$, it will not be joined into the region list. $\bar{c}(node)$ determines whether a node should be bisected further.
Discussion The advantages of hierarchical decomposition are at least three folds. Firstly, it well addresses the issue of object embedding and occlusion. In addition, the threshold $\tau_{1}$ is easier to set than the hyper-parameter k. The detected regions are the well reflection of the latent instance distribution in one image. The top-down decomposition reduces the bias accumulation caused by the bisecting and decreases the scale of the partition as the bisecting goes deeper. In the bisecting procedure, the partition could be achieved by NCut [28]. Nevertheless, the bisecting clustering is favored over NCut as the latter is much slower. Moreover, the decomposition by bisecting clustering is favored because it adapts well to the multi-scale semantic structure of an image, which will be revealed in the experiments.
@@
Method Descriptor Backbone Instance-160 Instance-335 INSTRE
mAP-50 mAP-100 mAP-all mAP-50 mAP-100 mAP-all mAP-50 mAP-100 mAP-all
@@
IV Feature Descriptor
The output from the repeated bisecting on the feature set $X$ is a hierarchy. Each node on the hierarchy is a subset of $X$ (except for the root node), which corresponds to a potential salient regions. In this section, we show how the features are extracted for these detected salient region, except for the dummy nodes on this hierarchy (discussed in Section III-D).
Given the location of the salient region is available, the feature extraction is as easy as performing RoI-pooling on a feature map from a network layer. To this end, we have multiple options. For the network, we can use the self-supervised ViT [9] based on which our salient regions are detected. Alternatively, the pre-trained ConvNet e.g., VGG [34] or ResNet [35] can be used to derive the feature. Moreover, instead of feeding the entire image into the network and pooling based on the region box, the detected region can be cropped out and fed into the network as one image to produce the feature. However, this way is computational inefficient. For this reason, we choose to feed the image into the pre-trained ResNet and extract the instance-level feature via RoI-pooling.
The complete pipeline of our hierarchical region detection and feature extraction is shown in Figure 3. On the first stage, the potential instance regions or sub-regions are detected via a top-down hierarchical decomposition over the feature set produced by backbone. Given the detection is an unsupervised procedure and is built upon class-agnostic self-supervised ViT and the feature is extracted from a ResNet. Since the detector is built upon class-agnostic self-supervised ViT the produced feature becomes a class-agnostic descriptor in the real sense. For this reason, our feature is called CLass Agnostic Instance-level Descriptor (CLAID) from now on. In contrast to traditional local feature, only 30 instance level regions per image are extracted.
@@ V-A Experiment Setup
The effectiveness of the proposed instance-level feature CLAID is evaluated on three benchmarks for instance search. They are Instance-160, Instance-335222https://drive.google.com/file/d/1ZxVUO3uInFqiFzXfuokQxalwMl1fEDgN
@@
and INSTRE [36]. Instance-160 contains 160 queries and 11,885 reference images. Instance-335 is an augmented dataset of Instance-160 with data collected from the GOT-10K [37], Youtube BoundingBoxes [38], and LaSOT [39] datasets. There are 335 instance queries and 40,914 reference images. For INSTRE dataset, there are 1,250 queries and 27,293 reference images. Compared to Instance-160 and Instance-335, INSTRE is rich of background variations and covers a wide range of object categories. In particular, there are 250 queries containing multiple instances. In order to verify the performance of CLAID in the multiple instance search task, the performance on this subset of queries (given as INSTRE-M) is further studied in the paper. The search quality is measured by mAP. Specifically, mAP-50, mAP-100, and mAP-all are reported for our method and state-of-the-art methods considered in the comparison. In addition, we also reported the instance localization accuracy on three datasets, which is measured by the commonly used Intersection Over Union (IoU). In order to study the exhaustiveness of our detector in discovering the latent instance regions in one image, we test it as an unsupervised object detection method. We study the recall of the correct localizations on COCO validation dataset [40] in comparison to state-of-the-art methods CutLER [10], TokenCut [12], and Deep Spectral [11].
Theoretically speaking, detector can be built upon any ViT networks. Due to its outstanding performance, ViT-Base (ViT-B/8) trained by DINO [9] is selected as the default backbone. For feature extraction backbone, there could be several options as well. ResNet-50  [35] is selected in our standard configuration follwoing the same setting in  [4] where feature is derived from Block4.0. Moreover, the search performance based on the other backbones such as ResNet-101 [35] and VGG-16 [34] is presented to validate the effectiveness of our detector. Both the detector and descriptor were implemented using PyTorch. The experiments were conducted on an Nvidia RTX 4090 GPU.
@@
TABLE II: Performance (mAP) on multi-instance search task
@@
TABLE III: Localization accuracy (mIoU) on Instance-160, Instance-335, and INSTRE. The IoU threshold is on 0.5 level

### v2

@@ III-B Decomposition Seed Selection
[…]
@@
$x_{b}=\underset{x}{\operatorname{arg\,min}}\,\sum A_{i{\cdot}}$ (6)
$x_{w}=\underset{x}{\operatorname{arg\,max}}\,\sum A_{i{\cdot}}$ (7)
So $x_{b}$ and $x_{w}$ are selected as two seeds for the bisecting. The rest of the features in $S$ are assigned to a cluster of $x_{b}$ or $x_{w}$ by checking whether they are closer to $x_{b}$ or $x_{w}$. The initialization scheme is adopted for the bisecting of all the nodes on the hierarchy. The difference between the clustering with seed selection and clustering with random initialization is illustrated in Figure 2.
@@ III-C Terminate Condition
(a) input image
(b) instance regions detected by CLAID
Figure 3: An illustration of detecting dense objects in different sizes and shapes by CLAID. Different instances are covered by different colors. For clarity, the instances correspond to the non-leaf nodes are not shown in figure (b).
Given a feature set (or a node) $S$ on the hierarchy, the total number of edges within the feature set $S$ is
$c(S)=\sum_{i,j\in S\&i<j}A_{ij}.$ (8)
@@
where $n=|S|$. Apparently, $\bar{c}(S)$ is in the range $[0,1]$. Since features inside $S$ correspond to a collection of fixed-size patches on the image, the higher $\bar{c}(S)$ is, the more these patches are semantically closer to each other. Compared to the parent node, $\bar{c}{(S)}$ from the child nodes is usually greater. As the decomposition continues, $\bar{c}{(S)}$ of the decomposed node becomes closer to 1.0. The bisecting over $S$ should not be carried out further when the connectivity within a feature set is sufficiently high. At this stage, the integral of these features may represent an inseparable instance region in the image. To this end, we can control the decomposition granularity by tuning the threshold $\tau_{1}$ on $\bar{c}(S)$. A larger threshold leads to finer decomposition granularity. In our implementation, when $\bar{c}(S)\geq{\tau_{1}}$, the bisecting on $S$ is not undertaken further.
Subjected to the granularity of ViT, the patch size in ViT cannot be arbitrarily small. The coverage of a ViT patch could be bigger than the size of a small instance. The fine-grained feature description of the small object becomes impossible. Therefore, the localization of such a small instance cannot be achieved by our method. Instead, a collection of dense small instances will be detected as a big virtual instance, namely an instance containing a group of dense small instances. This may not be meaningful for the semantic segmentation. Nevertheless, this is reasonable in the context of instance search. Since these small instances lose their individual details as objects, it is more realistic to treat them as a whole to be a subject in the image/instance search.
Figure 3 shows a typical example where dense objects are displayed in different sizes and shapes. As shown in the figure, our method is able to adapt to objects in different shapes and sizes. The dense small sunflowers in the distance are detected as a big instance (covered by yellow).
@@ III-D Condition on Dummy Nodes
(a) input image
(b) high energy regions
(c) dummy node (in blue)
(d) latent instances
Figure 4: The illustration of visual instances under the “dummy node” in a real scenario. A region (in blue) in Figure (c) is detected as “dummy node” based on Eqn. 10 after the 3rd bisecting. Two latent instances are detected when we further decompose the “dummy node” at the 5th bisecting. Figure (b) visualizes the high energy regions, i.e., $H$ (in red color) in the image.
Given a feature set (or a node) $S$ on the hierarchy, it does not necessarily correspond to a semantically salient region. It could be the mixture of objects and the background, or simply an empty background. We should not build a feature for such a region. Firstly, it is unnecessary to build a feature if the region is empty. Secondly, for the region that an instance is embedded in a big, empty background, the feature representation could be mixed with noise due to the dominant background. It would be more appropriate to build the feature when the empty background has been decomposed. We view such a region as a “dummy node” in the decomposition hierarchy. Inspired by [36], “dummy node” is judged by the strength of feature energy. In our method, we define $H=\arg\max\nolimits_{\forall x_{i}\in X}^{\theta}\left|x_{i}\right|^{1}$ as the set of features that exhibit high energy within the dataset $X$. Here, we set $\theta$ to the top-$30\%$ based on the relevant experimental results presented in [36]. We measure the overlapping ratio between $S$ and $H$
@@
When $\xi(S)$ is low, it basically indicates the region is over-dominated by semantically insignificant features. In our implementation, a node on the bisecting hierarchy is viewed as “dummy” if $\xi(S)\leq\tau_{2}$, where $\tau_{2}$ is empirically set. No feature will be built from it when a detected region is viewed as a “dummy node”. However, this condition does not prevent the feature set $S$ from further bisecting. Since $\xi(S)$ is a relative ratio, $\xi(S)$ may become higher when the bisecting reaches a more fine-grained level. Therefore, the direct/indirect subset of $S$ may become a salient region.
Figure 4 shows an example where “dummy node” is detected. After the 3rd bisecting, the region in blue (Figure 4(c)) is detected as “dummy node” for its relatively low intersection ratio with $H$. We do not simply ignore such a region. Instead, further bisecting is applied to the region. After another two times of bisecting on this node, two latent instances are successfully detected because their relative intersection ratio (i.e., $\xi(S)$) becomes much higher this time.
Figure 5: The framework for building instance-level features for instance search. There are two major components: an instance-level region detector and a feature descriptor. Given an image, a set of patch-level features is produced by the self-supervised backbone. The hierarchical decomposition is applied to the feature set. Each node on the hierarchy corresponds to a potential instance region or sub-region. The patch-level masks can be produced for the valid nodes on the hierarchy. The feature for each region is pooled from another network. Given an incoming query, a feature is extracted with the same backbone.
Input: Input image II
@@ 3 Initialize $\mathcal{L}\leftarrow X$ and $\mathcal{R}\leftarrow\emptyset$;
5  Pop nodenode from $\mathcal{L}$;
@@
7    $\mathcal{R}$.append(nodenode);
@@ 11    $\mathcal{L}\leftarrow\mathcal{L}\cup\text{getObject}(S_{b})\cup\text{getObject}(S_{w})$;
The entire process of instance-level detection is summarized in Algorithm 1. The hierarchical decomposition starts from the feature set $X$. The bisecting is repeatedly applied to $X$ and its subsets. The resulting subsets $S_{b}$ and $S_{w}$ do not necessarily correspond to two instance regions. Several instances from the same object category are likely partitioned into one subset. As a result, one more step $getObject(\cdot)$” (Algorithm 1 Line – 10) is undertaken to partition subsets $S_{b}$ and $S_{w}$ respectively into spatially connected smaller subsets. $getObject(\cdot)$” simply divides a subset into spatially connected regions by the eight-connectivity criterion, where tiny regions are ignored. The detection process results in a list of coherent candidate instances and their masks. Two subsidiary conditions are involved in the algorithm. Specifically, if a node is dummy according to $\xi(node)$, it will not be joined into the region list. $\bar{c}(node)$ determines whether a node should be bisected further.
Discussion Compared to CutLER [11] and TokenCut [13], we do not simply treat the two regions after the decomposition as either foreground or background. Instead, we believe that there are latent instances in both regions. Secondly, instead of setting a threshold on the number of decompositions explicitly [13, 11], we introduce the threshold $\tau_{1}$ on the average internal connectivity. It basically reflects the internal semantic similarities of image patches within a region. When the patches within a region show the same or similar semantics (reflected as a high $\bar{c}(S)$ value), it is believed to be an inseparable instance region and no need to be decomposed further. Moreover, compared to NCut [32] adopted by CutLER and TokenCut, the bisecting by k-sums [34] employed in our method is much efficient. As will be revealed in the experiments, it also adapts well to the multi-scale semantic structure of an image. Additionally, the introduction of “dummy node”, on the one hand, largely reduces the number of invalid regions that are dominated by semantically insignificant patches. On the other hand, it also allows us to discover small instances that are embedded in the big background.
IV Feature Descriptor
The output from the repeated bisecting on the feature set $X$ is a hierarchy. Each node on the hierarchy is a subset of $X$ (except for the root node), which corresponds to a potential salient region. In this section, we show how the features are extracted for these detected salient regions, except for the dummy nodes on this hierarchy (discussed in Section III-D).
Given the location of the salient region is available, the feature extraction is as easy as performing RoI-pooling on a feature map from a network layer. To this end, we have multiple options. For the network, we can use the self-supervised ViT [10], based on which our salient regions are detected. Alternatively, the pre-trained ConvNet, e.g., VGG [38] or ResNet [39] can be used to derive the feature. Moreover, instead of feeding the entire image into the network and pooling based on the region box, the detected region can be cropped out and fed into the network as one image to produce the feature. However, this way is computationally inefficient. For this reason, we choose to feed the image into the pre-trained ResNet and extract the instance-level feature via RoI-pooling.
The complete pipeline of our hierarchical region detection and feature extraction is shown in Figure 5. At the first stage, the potential instance regions or sub-regions are detected via a top-down hierarchical decomposition over the feature set produced by the backbone. Given that the detection is an unsupervised procedure and is built upon class-agnostic self-supervised ViT, and the feature is extracted from a ResNet. Since the detector is built upon class-agnostic self-supervised ViT, the produced feature becomes a class-agnostic descriptor in the real sense. For this reason, our feature is called CLass Agnostic Instance-level Descriptor (CLAID) from now on. In contrast to traditional local features, only 30 instance-level regions per image are extracted. For an instance query, the cropped query region is fed into the “Descriptor” block directly. Only one feature vector is produced for a query.
@@ IV Feature Descriptor
Method Descriptor Instance-160 Instance-335 INSTRE
backbone mAP-50 mAP-100 mAP-all mAP-50 mAP-100 mAP-all mAP-50 mAP-100 mAP-all
@@ V-A Experiment Setup
TABLE II: Summary over datasets used in the performance evaluation
Task Dataset #Queries #Images
Instance Search Instance-160 160 11,885
Instance-335 335 40,914
INSTRE 1,250 27,293
INSTRE-M 250 27,293
Image Retrieval Holidays 500 1,491
Oxford5K 55 5,062
Paris6K 55 6,412
The effectiveness of the proposed method CLAID is evaluated on both single instance search, multi-instance search, as well as image retrieval tasks. Three popular benchmarks in instance search are adopted. They are Instance-160, Instance-335222https://drive.google.com/file/d/1ZxVUO3uInFqiFzXfuokQxalwMl1fEDgN
@@
and INSTRE [40]. Instance-335 is an augmented dataset of Instance-160 with data collected from the GOT-10K [41], Youtube BoundingBoxes [42], and LaSOT [43] datasets. In INSTRE, there are 250 queries containing multiple instances. In order to verify the performance of CLAID in the multiple instance search task, the performance on this subset of queries (given as INSTRE-M) is further studied. In the image retrieval task, Holidays, Oxford5K, and Paris6K are adopted. The major information about these datasets are summarized in Tab. II. The search quality is measured by mAP. Specifically, mAP-50, mAP-100, and mAP-all are reported for instance search. mAP-all is reported for the image retrieval task.
In addition, we also report the instance localization accuracy on three datasets, which is measured by the commonly used Intersection Over Union (IoU). In order to study the exhaustiveness of our detector in discovering the latent instance regions in one image, we test it as an unsupervised object detection method. We study the recall of the correct localizations on COCO 2017 validation dataset with stuff and instance annotations [44] in comparison to state-of-the-art methods CutLER [11], TokenCut [13], Deep Spectral [12], and VoteCut [30].
Theoretically speaking, a detector can be built upon any ViT network. Due to its outstanding performance, ViT-Base (ViT-B/8) trained by DINO [10] is selected as the default backbone. For the feature extraction backbone, there could be several options as well. ResNet-50 [39] is selected in our standard configuration following the same setting in  [4] where the feature is derived from Block4.0. Moreover, the search performance based on the other backbones such as ResNet-101 [39] and VGG-16 [38] is presented to validate the effectiveness of our detector. Both the detector and descriptor were implemented using PyTorch. The experiments were conducted on an Nvidia RTX 4090 GPU.
@@
V-B Instance Search Performance
In the first experiment, the effectiveness of our method is studied in comparison to several representative methods in instance search. Among these methods, R-MAC [18], CAM-weight [45], BLCF [21], and BLCF-SalGAN [21] are deep global features that collapse all the features of an image into one vector for retrieval. In contrast, methods like DASR [4] and ours extract the instance-level feature from the candidate images and are able to localize the instance from the candidate images. In addition, we also treat the unsupervised object detectors CutLER [11], U2Seg [29], VoteCut [30], and Orth [31] as the instance region detectors. The same feature as ours is built for the regions detected by CutLER, U2Seg, VoteCut, and Orth, respectively. Moreover, we also report the results when the bisecting clustering in CLAID is replaced by NCut [32], while keeping other configurations unchanged. The instance search performance is shown in Table I.
As seen from the table, global features show poor performance on the instance search task. Object level features such as DASR, CutLER, U2Seg, VoteCut, Orth, and CLAID demonstrate considerably better performance. CLAID outperforms the rest of the methods on three benchmarks. Although methods such as CutLER, U2Seg, VoteCut, and Orth are also able to build class-agnostic instance-level features, their performance on instance search is inferior to CLAID. The relatively poor performance is mainly because they only detect the main objects in one image. Typically, the cut in a chain in CutLER causes the loss of potential instance regions. Similar comments apply to U2Seg and VoteCut. Compared to CLAID (NCut), CLAID performs considerably better. According to our observation, the results from bisecting clustering better reflect the multi-scale semantic structure than the partitions produced by NCut. In addition, we also test CLAID with different backbones at its feature extraction stage. As shown in the table, our method shows considerably better performance when it is supported by the same backbone as the competing methods (e.g., DASR and CutLER).
Figure 6 further showcases the instance search results from CLAID. More instance search results can be found from the Appendix. Meanwhile, the performance on other pre-trained ViT networks such as DINOv2 [46], SigLIP [47], and ViT pretrained by  [33] is also reported in the Appendix.
In order to study the behavior of our method on a multi-instance search task, the performance on 250 queries from INSTRE is particularly evaluated. For these queries, there is more than one object present in a query. As shown in Table III, most of the methods show relatively better performance on this task than that of single instance search (reported in Table I). The search quality from all global features is considerably better. This is mainly because the large coverage of the instance query over the image makes it close to a conventional content-based image search task. The mixture of features from different instances becomes helpful in this context. Our method, along with recent class-agnostic detectors such as CutLER, U2Seg, VoteCut, and Orth, shows the best performance. All these methods are able to detect the major instance regions in one image, which are usually an integral of multiple instances.
TABLE III: Performance (mAP) on multi-instance search task
@@ V-B Instance Search Performance
The localization accuracy of the target instances on the candidate images is reported in Table IV. Since the localization is impossible to fulfill for the global features, only the performance from the level features is reported. The patch-level region masks from our method, CutLER, U2Seg, VoteCut, and Orth are transformed to bounding boxes when we calculate the mIoU score. As shown in Table IV, CLAID shows the secondary localization accuracy to VoteCut, which is the result of voting from several CutLER-like detectors based on six ViT backbones. The high localization accuracy is at the cost of relatively low detection efficiency (as shown in Table V).
V-C The Performance of Feature Detector
Figure 7: Recall at different IoU thresholds for seven methods on COCO 20k.
Since the feature detector is the critical component in our pipeline, its effectiveness is further studied from two perspectives, namely the detection speed and the exhaustiveness in detecting latent instance regions from an image. The processing speed is compared to unsupervised object detectors such as Deep Spectral [12], TokenCut [13], and CutLER [11]. All of them are built on self-supervised ViT. For this reason, we only evaluate the time cost of processing one image. The speed efficiency is evaluated on CPU with a single thread.
[…]
To verify the ability of our detector in detecting the latent instance regions in an arbitrary image, we further study the performance of CLAID on COCO 20k segmentation validation dataset. In this evaluation, CLAID is treated as an unsupervised instance detection method. The exhaustiveness is measured by the recall rate of the object instances in different categories that have been annotated in the dataset. The recall is evaluated with respect to IoU score. Six detectors Deep Spectral [12], TokenCut [13], CutLER [11], U2Seg[29], Orth[31] and VoteCut [30] are considered in the comparison. The detection performance is shown in Figure 7. As seen from the figure, CLAID outperforms these detectors by a large margin. On the one hand, it largely owes to the hierarchical decomposition without the depth constraint. It allows the exhaustive detection of the latent instances. On the other hand, the introduction of “dummy node” filters out the detected invalid instance regions.
TABLE IV: Localization accuracy (mIoU) on Instance-160, Instance-335, and INSTRE.